OpenStudy (anonymous):

Find the slope of the line passing through (5, 5) and (-4, 5)

OpenStudy (anonymous):

see in general ,the equation of the line passsing through two points (x1,y1) and (x2,y2) with slope m is: y-y1=m(x-x1) where m is (y2-y1)/(x2-x1)

OpenStudy (anonymous):

hope u get it..........

OpenStudy (anonymous):

find slope is about find tan of the angle of the line with horizontal line, so you can figure that by formula: slope = (y2-y1)/(x2-x1)